What is the mass percentage of 125g of fructose dissolved in 375mL of water?


1
Expert's answer
2022-03-22T04:22:04-0400

% mm = m(fructose) / m(solution) × 100% = 125 g / (125 g + 375 g) × 100% = 125 g / 500 g × 100% = 25%